Calculate the molar enthalpy of vaporization of a liquid which boils at 120°C.
1
Expert's answer
2019-10-25T06:26:01-0400

Using Clausius-Clayperon equation "lnP=-\frac{∆H}{RT}"

"P=10^5Pa"

"∆H=37.624\frac{KJ}{mol}"